Find the area bounded by the curves y=4/x^2, x-2y+7 and x=4.


1
Expert's answer
2021-04-25T07:36:31-0400

At the intersection of the curves  

"y=\frac4{x^2}" and "x-2y+7=0" we have "\frac 4{x^2} = \frac {x+7}2"

0r "8=x^3+7x" or "x^3+7x-8=0 \u2026(i)"

Which is a cubic equitation with x=1 as a root.

We thus use synthetic division to find other roots (if any).

"\begin {array}{c| rrr}&1&1&0&-8\\-&&1&1&8\\\hline\\&1&1&2&0\\\end{array}"


"\u21d2x^3+7x-8= (x-1)(x^2+x-8)=0"

The function   "x^2+x+8" has no real roots suggesting that the curves  "y= \frac 4{x^2}" and

"x+2y+7=0" intersect at a point where x=1.

"⇒area= ∫_1^4(\frac {x+7}2 - \frac 4{x^2})dx"

"=[\frac{ x^2}{4}+ \frac 72 x+\frac 4x]_1^{4}"

"=9-\frac {31}4"

"=11.25 \, square\, units\, of \, length"